OpenStudy (anonymous):

has anyone done the imploding can experiment where you calculate efficiency!

OpenStudy (anonymous):

@surry99 Can you help us?

OpenStudy (surry99):

sure....set up the problem

OpenStudy (anonymous):

the efficiency n = w/Q where W is work and Q is heat input. derive an equation for n, and in the simplest form all quantities should be known. we know the volume of the can, the atmospheric pressure, latent heat, and mass of steam

OpenStudy (anonymous):

\[n=(P \Delta V)/L(m)\] or is there more to it?

OpenStudy (surry99):

ok it looks like you... 1) filled a can with water 2) heated it until it became steam 3) let it out of the can 4) the can was then crushed by the atmospheric pressure correct?

OpenStudy (anonymous):

no... after your step two... we grabed it with tongs, quickly turned it upside down and submerged it into ice water in which in imploded. Then filled it with water to measure the new volume

OpenStudy (anonymous):

Hold on i'll post you a photo that might help you better understand

OpenStudy (surry99):

oh yes that makes sense...then your formula is correct, Work = Patm(deltaV) Q = mL

OpenStudy (anonymous):

We were questioning it because we are getting numbers that would give us like 50% efficiency. Our lab instructor told us an engine of a car has like a 20% efficiency so we thought the can should be less. But the can shouldn;t be less because there are less outside sources working against it, correct?

OpenStudy (anonymous):

do you know the units that we need to plug into that equation?

OpenStudy (surry99):

No your results sound reasonable and car engines are not very efficient for a variety of reasons. Good question about units...does an efficiency have units?

OpenStudy (anonymous):

We can't find any, when we look at example problems in our text book they have commas next to them... thats why I suggested maybe putting the answer in percentages.

OpenStudy (surry99):

Your efficiency should have no units. The units for work on the top should be the same as the units for heat on the bottom. What units did you use for: P V M L

OpenStudy (anonymous):

efficiency is a ratio between heat going in and heat going out. An equation i found in my book says that efficiency = 1 - Qinput/Qoutput.......which is why i am questioning our initial equation. if our initial equation says the numerator is W = PdeltaV, and the book equation says 1 - Qin / Qout

OpenStudy (surry99):

That is correct ...it is for a Carnot engine....but did you in anyway determine Qin and Q out. My guess is you know: P in atmospheres V was milliliters m was in grams L was in Joules /gram You need to let me know if these are correct.

OpenStudy (anonymous):

I think the book said Latent was kj... i'm trying to find the page...not 100% on that

OpenStudy (anonymous):

those are correct, which i then further converted into P in N/msquared V in m cubed Latent in J/kg Mass in kg therefor i think it all crosses out to be in Joules/Joules....but that number is VERY small. I did a few calculations and got less than 10% efficient

OpenStudy (surry99):

ok lets summarize: 1) My mistake for indicating 50% was reasonable...I read it fast and thought it said 5%. 2) Physicsgirls units indeed cancel out as they should...excellent work 3) n= 1 - qin/qout is derived from n=w/q....same equation 4) less than 10% looks better

OpenStudy (anonymous):

Yeah sorry about the 50% she went on and redid the problem and forget to mention it to me :-P

OpenStudy (anonymous):

so then, how do we find Qin? we only have information for PdeltaV..... Qin would be L*m....but do we know that information?

OpenStudy (surry99):

PdeltaV is correct for the work...remember the equations are the same and you measured P and deltaV

OpenStudy (anonymous):

i guess a better question is: how can n = w/q be equal to n = 1- Qin/Qout when W = Qin - Qout

OpenStudy (anonymous):

does that mean PdeltaV = Qin - Qout?

OpenStudy (surry99):

hang on...

OpenStudy (surry99):

yes

OpenStudy (surry99):

Gotta run...you two ok?

OpenStudy (anonymous):

yes real quick, if our effency equation comes up with 10 percent, and the carnot equation comes up with 26% is that a reasonable comparison

OpenStudy (surry99):

yes that is...the Carnot efficiency is the best you can possible do if you try to convert heat from a high temperature source to do work because no matter what you will have to release some heat at a lower temperature. That is it, it is not possible to convert all the heat to do work...some heat must always be rejected.

OpenStudy (anonymous):

thank you so much!

OpenStudy (surry99):

You are both very welcome. Goodnight to you both.